# On Experiments

### v1

@@ 5 Experiments
We have assumed that the decision maker represents their uncertainty in $\Theta$ via a probability distribution $P$. The key question is which $P$. Under the assumption that the decision maker aims to minimize their expected loss under $P$, this question is equivalent to asking which admissible action $a_{P}$ the decision maker should use. To ascertain $P$, the decision maker is guided by experiments.
In its simplest form, an experiment is a function $e:\Theta\rightarrow\mathbb{P}(\mathcal{Z})$, that assigns to each value $\theta$ a probability distribution over the outcomes of the experiment. By linearity, this function can be extended to a linear function $(\mathbb{R}^{\Theta})^{*}\rightarrow(\mathbb{R}^{\mathcal{Z}})^{*}$. It is these more general objects that will take the focus of our study.
@@ 5.1 Transitions and their Algebra
While abstract in appearance, we remark that when $X$ and $Y$ are finite, a transition is nothing more than a matrix. In general, a transition is an integral operator. Denote the set of all transitions from $X$ to $Y$ by $\mathbb{T}(X,Y)$. We call a transition Markov if $T(\mathbb{P}(X))\subseteq\mathbb{P}(Y)$, ie it maps distributions over $X$ to distributions over $Y$. Denote the set of all Markov transitions from $X$ to $Y$ by $\mathbb{M}(X,Y)$. When $X$ and $Y$ are finite, Markov transitions are represented by column stochastic matrices. The distribution,
@@
For transitions $f\in\mathbb{T}(X,Y)$ and $g\in\mathbb{T}(Y,Z)$ we can define $g\circ f\in\mathbb{T}(X,Z)$ by usual function composition. If $f_{i}\in\mathbb{T}(X_{i},Y_{i})$, $i\in[1;k]$, are transitions then denote,
@@
with $\otimes_{i=1}^{k}f_{i}(x)=f_{1}(x_{1})\otimes\dots\otimes f_{k}(x_{k})$, where $\times$ denotes the Cartesian product and $\otimes$ denotes products of duals. Transitions can also be replicated. For any transition $f\in\mathbb{T}(X,Y)$ we denote the replicated transition $f_{n}\in\mathbb{T}(X,Y^{n}),\ n\in\{1,2,\dots\}$, with,
@@
The above can be understood as definining a particular category, with objects as finite sets and arrows as transitions. Restricting to Markov transitions yields a subcategory, sometimes called the category of stochastic relations. It is this category that will be of our prime interest.
@@ 5.2 Comparing Experiments
An experiment is a Markov transition $e\in\mathbb{M}(\Theta,\mathcal{Z})$. We call $\mathcal{Z}$ the observation space of the experiment. We assume that $\mathcal{Z}$ is finite. The distribution $e(\theta)$ summarizes the decision maker’s uncertainty in the observation when $\theta$ is the value of the unknown. Allowing $\mathcal{Z}$ to vary yields a category $\mathbb{M}(\Theta,-)$, of elements under $\Theta$. Objects of this category are experiments with arrows given by commutative triangles,
@@
The Markov transition $f$ can be understood as adding noise to the experiment $e$, yielding the corrupted experiment $e^{\prime}$. After observing the results of an experiment, the decision maker is tasked with choosing a suitable action. They do this via a decision rule.
A decision rule is a Markov transition $d\in\mathbb{M}(\mathcal{Z},A)$. The distribution $d(z)$ summarizes the decision makers uncertainty in which action to choose, given an observation $z\in\mathcal{Z}$. We define the risk,
@@
The risk measures the quality of the final action chosen by the decision maker when they use the decision rule $d$, after performing experiment $e$, assuming $\theta$ is the true value of the unknown. The risk does not provide a single number for the comparison of experiments, rather it provides an entire risk profile. To compare risks directly, the decision maker can use the Bayesian or max risks defined as,
@@
respectively. The Bayesian risk is more appropriate if the decision maker has some intuition about $\theta$, given in the form of a prior probability distribution $\pi$. The max risk is more appropriate if the decision maker has no prior knowledge concerning $\theta$. These quantities allow the decision maker to compare the usefulness of experiment, decision rule pairs. To compare experiments directly, we assume the decision maker uses the best decision rule. Define the minimum Bayesian risk and minimax risk as,
@@
$\underline{{\mathrm{Risk}}}_{L}(e)=\sup_{\pi\in\mathbb{P}(\Theta)}\underline{{\mathrm{Risk}}}_{L}^{\pi}(e).$
The proof is a simple application of the minimax theorem [20]. In light of this theorem, we focus on Bayesian risks for the remainder. All results have a minimax equivalent.
@@ 5.3 Risk as Loss
Abstracting away the details of experiment, what is important to the decision maker is the quality of the composed transition $d\circ e\in\mathbb{M}(\Theta,A)$. One could call such a transition a strategy. Risk is a function,
@@ 5.4 Bias-Variance Decomposition for Risks
When working with canonical losses, the risk admits a simple, understandable decomposition. For canonical losses we have,
@@
$=\mathbb{E}_{Z}-\left<\delta_{\theta},d(Z)\right>+\Psi(d(Z))$
$=-\left<\delta_{\theta},\bar{d}\right>+\Psi(\bar{d})+\mathbb{E}_{Z}\Psi(d(Z))-\Psi(\bar{d})$
@@ 5.5 Admissible and Bayesian decision rules
The optimal decision rule will in general depend on the decision makers prior knowledge about the unknown. Even without this knowledge, the decision maker can remove rules that are obviously not optimal.
@@
Intuitively, a decision rule is admissible if it is not obviously worse than some other decision rule. If the decision maker has prior $\pi$, they can minimize the Bayesian risk by using a Bayesian decision rule.
@@
Let $e$ be an experiment and $\pi$ a prior. A decision rule $d^{*}$ is Bayes for $(\pi,e)$ if,
@@
The above theorem says that Bayesian algorithms provide all rules that a sensible decision maker should use. Picking a particular admissible algorithm is equivalent to picking a prior $\pi$ and minimizing the Bayesian risk against that prior. While statistically speaking, admissible algorithms afford no obvious improvements, they may be hard to implement. Our language as is does not take this into account. The study of inadmissible algorithms and their risks is therefore a worthwhile endeavor.
Bayes optimal algorithms admit a simple representation. No paper on decision theory would be complete without reference to Bayes’ rule. In our language Bayes’ rule provides means to reverse an experiment. Let $e(\pi)$ be the marginal distribution over the observation space, and $e^{\dagger}\in\mathbb{M}(\mathcal{Z},\Theta)$ be the induced Markov transition of unknowns given observations obtained via Bayes rule. Then,
@@
$d(z)$ proceeds by first inferring which $\theta$ is true, before acting optimally. Denote by $\bullet$ the canonical one element set. To include prior distributions in our language, we can work with the category of elements under $\bullet$, denoted by $\mathbb{M}(\bullet,-)$. Objects of this category are precisely prior distributions defined over the relevant set. Arrows are commutative triangles that ensure the priors match up, ie an arrow,
@@ 6 When is One Experiment Always Better than Another?
Let $e$ and $e^{\prime}$ be experiments. Suppose that due to constraints, the decision maker can only perform one of these two experiments. The decision maker can compare the Bayes or minimax risks of the two experiments, however this involves performing a calculation. Furthermore, if the loss function of interest changes, then the ordering of the experiments might change. We seek qualitative results concerning when $e$ is always better than $e^{\prime}$ no matter what the loss or prior distribution.
@@
Intuitively, $e\mid e^{\prime}$ if $e^{\prime}$ is $e$ with extra noise $f$. We make this intuition precise with theorem 22. For an experiment $e$, let $\mathbb{M}_{e}(\Theta,A)$ be the set of transitions from $\Theta$ to $A$ that $e$ divides.
@@
As $e^{\prime}\in\mathbb{M}_{e^{\prime}}(\Theta,\mathcal{Z}^{\prime})$, this implies there exists a $f$ with $f\circ e=e^{\prime}$.
@@ 6.1 The Blackwell-Sherman-Stein Theorem and Sufficiency
We prove the forward implication, called the data processing theorem. The proof of the converse will come later, as a simple corollary of the randomization theorem.
@@
We say $e$ and $e^{\prime}$ are equivalent experiments (written $e\cong e^{\prime}$) if both $e\mid e^{\prime}$ and $e^{\prime}\mid e$. Equivalent experiments have equivalent risks. A key notion in statistics is that of sufficiency. A sufficient statistic is a function of the observation that loses none of the information contained in $e$. Identifying and exploiting sufficient statistics allows the decision maker to compress the information contained in the observation, without losing information.
@@ 6.1.1 Sufficiency and Bayes’ Rule
To see this, remember that decision rule that minimizes the Bayes risk is precisely that which infers $\theta$ via $e^{\dagger}$ and then acts optimally. By the Blackwell-Sherman-Stein theorem, $e^{\dagger}$ is sufficient for $e$.
@@ 6.2 Most and Least Informative Experiments
For any set $\Theta$ of unknowns there is a most informative and a least informative experiment. Recall the identity function $\mathrm{id}_{\Theta}$, $\mathrm{id}_{\Theta}(\theta)=\theta$. For any experiment $e$, we have $e\circ\mathrm{id}_{\Theta}=e$. Therefore, $\mathrm{id}_{\Theta}$ divides any experiment. Intuitively, $\mathrm{id}_{\Theta}$ provides the decision maker the exact value of $\theta$. This experiment has risk,
@@
For any set $X$, define the terminal transition $\bullet_{X}\in\mathbb{M}(X,\{1\})$ with $\bullet_{X}(x)=1$ for all $X$. Intuitively this transition throws away all information about $X$. Much like the identity transition divides every experiment, the terminal transition is divided by every experiment. For all experiments $e$, $\bullet_{\Theta}=\bullet_{\mathcal{Z}}\circ e$. This experiment has risk,
@@ 6.3 The General Data Processing Theorem
Key to the proof of the data processing theorem is the insight that if $e\mid e^{\prime}$ then,
@@
Intuitively, this means more decision rules are available to the decision maker after performing the experiment $e$ than $e^{\prime}$, as one could always first inject noise and then proceed as if experiment $e^{\prime}$ had been performed. This suggests another means to construct quantities that satisfy data processing theorems.
@@
Remarkably, the proof of theorem 24 makes no explicit reference to expected risks, transitions, or even probability distributions. Digging deeper, one can understand data processing theorems as functors. Define the category $\left(\mathbb{R},\leq\right)$, with objects real numbers and arrows $x\rightarrow y$ if and only if $x\leq y$. Similarly consider the category of finite sets with arrows Markov transitions. Fix a set of unknowns $\Theta$ and consider the experiments on $\Theta$, ie elements of $\mathbb{M}(\Theta,-)$. This is naturally a category, sometimes called a slice category or the category of elements under $\Theta$, with arrows given by commutative triangles linking experiments via noise,
@@
Then the construction from theorem 24 can be understood as a functor from $\mathbb{M}(\Theta,-)$ to $\left(\mathbb{R},\leq\right)$.
@@ 6.4 Understanding the Blackwell-Sherman-Stein Theorem as an Equivalence of Categories
Then the Blackwell-Sherman-Stein theorem then states that these categories are equivalent, ${\mathrm{Risk}}(\Theta)\cong\mathbb{M}(\Theta,-)$.
@@ 6.5 An Example: Binary Experiments and the Variational Divergence
Let $\Theta=A=\{-1,1\}$ with $L(\theta,a)=[\![\theta\neq a]\!]$. In words, there are two unknowns, and all the loss function cares about is that we accurately identify which of the two it is. Let $e$ be an experiment, and $\pi$ the uniform prior on $\Theta$. Then,
@@
which is the total variational distance between e(1)e(-1) and $e(1)$, a commonly used metric between probability distributions. The data processing theorem recovers the well known fact that,
@@
for all distrbutions $P,Q\in\mathbb{P}(\mathcal{Z})$ and all Markov kernels $f\in\mathbb{M}(\mathcal{Z},\mathcal{Z}^{\prime})$.
@@ 6.6 Example: General Experiments and the Mutual Information
Let $\phi:[0,\infty)\rightarrow\mathbb{R}$ be a convex function with $f(1)=0$. Then the $f$-divergence between distributions $P,Q\in\mathbb{P}(\mathcal{Z})$ is given by,
$D_{f}(P,Q)=\mathbb{E}_{P}f(\frac{dQ}{dP}),$
@@
for some $\phi$. Similarly for every $\phi$-divergence, there is a loss and prior that makes the above hold. We refer the reader to [8] for the details of this correspondence. Our general data processing theorem then recovers the fact that,
@@
and objects called multi-$\phi$-divergences. We refer the reader to [22] for the details. Our general data processing theorem once again recovers the data processing theorem for multi-$\phi$-divergences.

### v2

@@ 5 Experiments
We have assumed that the decision maker represents their uncertainty in $\Theta$ through a probability distribution $P$. The key question is which $P$. Under the assumption that the decision maker aims to minimize their expected loss under $P$, this question is equivalent to asking which admissible action $a_{P}$ the decision maker should use. To determine $P$, the decision maker is guided by experiments.
In its simplest form, an experiment is a function $e:\Theta\rightarrow\mathbb{P}(\mathcal{Z})$, which assigns to each value $\theta$ a probability distribution on the results of the experiment. By linearity, this function can be extended to a linear function $(\mathbb{R}^{\Theta})^{*}\rightarrow(\mathbb{R}^{\mathcal{Z}})^{*}$. These more general objects will be the focus of our study.
@@ 5.1 Transitions and their Algebra
Although abstract in appearance, we observe that when $X$ and $Y$ are finite, a transition is nothing more than a matrix. In general, a transition is an integral operator. Denote the set of all the transitions from $X$ to $Y$ by $\mathbb{T}(X,Y)$. We call a transition Markov if $T(\mathbb{P}(X))\subseteq\mathbb{P}(Y)$. Denote the set of all Markov transitions from $X$ to $Y$ by $\mathbb{M}(X,Y)$. When $X$ and $Y$ are finite, Markov transitions are represented by column stochastic matrices. The distribution,
@@
For transitions $f\in\mathbb{T}(X,Y)$ and $g\in\mathbb{T}(Y,Z)$ we can define $g\circ f\in\mathbb{T}(X,Z)$ by the composition of functions. If $f_{i}\in\mathbb{T}(X_{i},Y_{i})$, $i\in[1;k]$, are transitions then denote,
@@
with $\otimes_{i=1}^{k}f_{i}(x)=f_{1}(x_{1})\otimes\dots\otimes f_{k}(x_{k})$, where $\times$ denotes the Cartesian product and $\otimes$ denotes dual products. Transitions can also be replicated. For any transition $f\in\mathbb{T}(X,Y)$ we denote the replicated transition $f_{n}\in\mathbb{T}(X,Y^{n}),\ n\in\{1,2,\dots\}$, with,
@@
The above can be understood as defining a particular category, with objects as finite sets and arrows as transitions. Restricting to Markov transitions yields a subcategory, sometimes called the category of stochastic relations. This category will be of our primary interest.
@@ 5.2 Comparing Experiments
An experiment is a Markov transition $e\in\mathbb{M}(\Theta,\mathcal{Z})$. We call $\mathcal{Z}$ the observation space of the experiment. We assume that $\mathcal{Z}$ is finite. The distribution $e(\theta)$ summarizes the uncertainty of the decision maker in the observation when $\theta$ is the unknown value. Leaving $\mathcal{Z}$ to vary yields a category $\mathbb{M}(\Theta,-)$, of elements under $\Theta$. Objects of this category are experiments with arrows given by commutative triangles,
@@
The Markov transition $f$ can be understood as adding noise to the experiment $e$, which results in the corrupted experiment $e^{\prime}$. After observing the results of an experiment, the decision maker is tasked with choosing a suitable action. They do this through decision rules.
A decision rule is a Markov transition $d\in\mathbb{M}(\mathcal{Z},A)$. The distribution $d(z)$ summarizes the uncertainty of decision maker in which action to choose, given an observation $z\in\mathcal{Z}$. We define the risk,
@@
The risk measures the quality of the final action chosen by the decision maker when using the decision rule $d$, after performing the experiment $e$, assuming $\theta$ is the true value of the unknown. The risk does not provide a single number for the comparison of experiments; rather, it provides an entire risk profile. To compare risks directly, the decision maker can use Bayesian or max risks defined as
@@
respectively. Bayesian risk is more appropriate if the decision maker has some intuition about $\theta$, given in the form of a prior probability distribution $\pi$. The maximum risk is more appropriate if the decision maker has no prior knowledge of $\theta$. These quantities allow the decision maker to compare the usefulness of experiment, decision rule pairs. To compare experiments directly, we assume that the decision maker uses the best decision rule. Define the minimum Bayesian risk and minimax risk as,
@@
$\underline{{\mathrm{Risk}}}_{L}(e)=\max_{\pi\in\mathbb{P}(\Theta)}\underline{{\mathrm{Risk}}}_{L}^{\pi}(e).$
The proof is a simple application of the minimax theorem [11]. In light of this theorem, we focus on Bayesian risks for the remainder.
@@ 5.3 Risk as Loss
Leaving out the details of the experiment, what is important to the decision maker is the quality of the composed transition $d\circ e\in\mathbb{M}(\Theta,A)$. One could call such a transition a strategy. Risk is a function,
@@ 5.4 Bias-Variance Decomposition for Risks
When working with canonical losses, the risk admits a simple and understandable decomposition. For canonical losses, we have,
@@
$=\mathbb{E}_{Z}\left<\delta_{\theta},d(Z)\right>+\Psi(d(Z))$
$=\left<\delta_{\theta},\bar{d}\right>+\Psi(\bar{d})+\mathbb{E}_{Z}\Psi(d(Z))-\Psi(\bar{d})$
@@ 5.5 Admissible and Bayesian decision rules
The optimal decision rule will, in general, depend on the prior knowledge of the decision maker about the unknown. Even without this knowledge, the decision maker can remove rules that are obviously not optimal.
@@
A decision rule is admissible if it is not obviously worse than some other decision rule. If the decision maker has prior knowledge $\pi$, they can minimize the Bayesian risk using a Bayesian decision rule.
@@
Let $e$ be an experiment and $\pi$ be a prior. A decision rule $d^{*}$ is Bayes for $(\pi,e)$ if,
@@
The above theorem says that Bayesian algorithms provide all the rules that a sensible decision maker should use. Picking a particular admissible algorithm is equivalent to picking a prior $\pi$ and minimizing the Bayesian risk against that prior. Although statistically speaking, admissible algorithms afford no obvious improvements, they may be hard to implement. Our language does not take this into account. The study of inadmissible algorithms and their risks is, therefore, a worthwhile endeavor.
Bayes optimal algorithms admit a simple representation. No paper on decision theory would be complete without reference to Bayes’ rule. In our language, Bayes’ rule provides means to reverse an experiment. Let $e(\pi)$ be the marginal distribution over the observation space, and $e^{\dagger}\in\mathbb{M}(\mathcal{Z},\Theta)$ be the induced Markov transition of unknowns given observations obtained via Bayes’ rule. Then,
@@
$d(z)$ proceeds first by inferring which $\theta$ is true, before acting optimally. Denote by $\bullet$ the canonical set of one element. To include prior distributions in our language, we can work with the category of elements under $\bullet$, denoted by $\mathbb{M}(\bullet,-)$. Objects of this category are precisely prior distributions defined over the relevant set. Arrows are commutative triangles that ensure the priors match up, ie an arrow,
@@ 6 When is One Experiment Always Better than Another?
Let $e$ and $e^{\prime}$ be experiments. Suppose that due to constraints, the decision maker can only perform one of these two experiments. The decision maker can compare the Bayes or minimax risks of the two experiments, however, this involves performing a calculation. Furthermore, if the loss function of interest changes, then the ordering of the experiments might change. We seek qualitative results in relation to when $e$ is always better than $e^{\prime}$ no matter what the loss or prior distribution.
@@
$e\mid e^{\prime}$ if $e^{\prime}$ is $e$ with extra noise $f$. We make this intuition precise with theorem 22. For an experiment $e$, let $\mathbb{M}_{e}(\Theta,A)$ be the set of transitions from $\Theta$ to $A$ that $e$ divides.
@@
As $e^{\prime}\in\mathbb{M}_{e^{\prime}}(\Theta,\mathcal{Z}^{\prime})$, this implies that there exists a $f$ with $f\circ e=e^{\prime}$.
@@ 6.1 The Blackwell-Sherman-Stein Theorem and Sufficiency
We prove the forward implication, called the data processing theorem. The proof of the converse will come later as a simple corollary of the randomization theorem.
@@
We say $e$ and $e^{\prime}$ are equivalent experiments (written $e\cong e^{\prime}$) if both $e\mid e^{\prime}$ and $e^{\prime}\mid e$. Equivalent experiments have equivalent risks. A key notion in statistics is that of sufficiency. A sufficient statistic is a function of the observation that loses none of the information contained in $e$. Identifying and exploiting sufficient statistics allows the decision maker to compress the information contained in the observation without losing information.
@@ 6.1.1 Sufficiency and Bayes’ Rule
To see this, remember that the decision rule that minimizes the Bayes risk is precisely that which infers $\theta$ through $e^{\dagger}$ and then acts optimally.
@@ 6.2 Most and Least Informative Experiments
For any set $\Theta$ of unknowns, there is a most informative and a least informative experiment. Recall the identity function $\mathrm{id}_{\Theta}$, $\mathrm{id}_{\Theta}(\theta)=\theta$. For any experiment $e$, we have $e\circ\mathrm{id}_{\Theta}=e$. Therefore, $\mathrm{id}_{\Theta}$ divides any experiment. $\mathrm{id}_{\Theta}$ provides the decision maker the exact value of $\theta$. This experiment has risk,
@@
For any set $X$, define the terminal transition $\bullet_{X}\in\mathbb{M}(X,\{1\})$ with $\bullet_{X}(x)=1$ for all $X$. This transition eliminates all information about $X$. Much like the identity transition divides every experiment, the terminal transition is divided by every experiment. For all experiments $e$, $\bullet_{\Theta}=\bullet_{\mathcal{Z}}\circ e$. This experiment has risk,
@@ 6.3 The General Data Processing Theorem
The key to the proof of the data processing theorem is the insight that if $e\mid e^{\prime}$ then,
@@
This means that the decision maker has more decision rules after performing the experiment $e$ than $e^{\prime}$, since one could always first inject noise and then proceed as if experiment $e^{\prime}$ had been performed. This suggests another means to construct quantities that satisfy the data processing theorem.
@@
Remarkably, the proof of theorem 24 makes no explicit reference to expected risks, transitions, or even probability distributions. One can understand data processing theorems as functors. Define the category $\left(\mathbb{R},\leq\right)$, with objects real numbers and arrows $x\rightarrow y$ if and only if $x\leq y$. Consider the category of finite sets with arrows Markov transitions. Fix a set of unknowns $\Theta$ and consider experiments on $\Theta$, elements of $\mathbb{M}(\Theta,-)$. This is naturally a category, sometimes called a slice category or the category of elements under $\Theta$, with arrows given by commutative triangles linking experiments via noise,
@@
Then the construction of Theorem 24 can be understood as a functor from $\mathbb{M}(\Theta,-)$ to $\left(\mathbb{R},\leq\right)$.
@@ 6.4 Understanding the Blackwell-Sherman-Stein Theorem as an Equivalence of Categories
Then the Blackwell-Sherman-Stein theorem states that these categories are equivalent ${\mathrm{Risk}}(\Theta)\cong\mathbb{M}(\Theta,-)$.
@@ 6.5 An Example: Binary Experiments and the Variational Divergence
Let $\Theta=A=\{-1,1\}$ with $L(\theta,a)=[\![\theta\neq a]\!]$. In words, there are two unknowns, and all the loss function cares about is that we accurately identify which of the two it is. Let $e$ be an experiment and $\pi$ the uniform prior on $\Theta$. Then,
@@
which is the total variational distance between e(1)e(-1) and $e(1)$, a commonly used metric between probability distributions. The data processing theorem recovers the well-known fact that,
@@
for all distributions $P,Q\in\mathbb{P}(\mathcal{Z})$ and all Markov kernels $f\in\mathbb{M}(\mathcal{Z},\mathcal{Z}^{\prime})$.
@@ 6.6 Example: General Experiments and the Mutual Information
Let $\phi:[0,\infty)\rightarrow\mathbb{R}$ be a convex function with $\phi(1)=0$. Then the $\phi$-divergence between the distributions $P,Q\in\mathbb{P}(\mathcal{Z})$ is given by,
$D_{\phi}(P,Q)=\mathbb{E}_{P}\phi\left(\frac{dQ}{dP}\right),$
@@
for some $\phi$. Similarly, for every $\phi$-divergence, there is a loss and prior that makes the above hold. We refer the reader to [16] for the details of this correspondence. Our general data processing theorem then recovers the fact,
@@
and objects called multi-$\phi$-divergences. We refer the reader to [8]. Our general data processing theorem once again recovers the data processing theorem for multi-$\phi$-divergences.